Circular No. 8766

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006jf–2006ld J. Frieman, on behalf of the Sloan Digital Sky Survey II collaboration,

reports the discovery of 48 new supernovae (2006jf–2006kd and 2006kg– 2006lc; magnitude range g = 19.8–23.3) on images taken during Sept. 11– Oct. 21 (details on CBETs 680 and 688). All appear to be apparent or possible type-Ia supernovae except for 2006jo and 2006jx (apparent or possible type Ib); 2006kg, 2006kh, 2006kn, and 2006kv (apparent or possible type II); 2006jl (IIn); and 2006lc (probable type-Ic).

In addition, three additional apparent supernovae have been discovered on unfiltered CCD images: 2006ke and 2006kf by J. Schwehr, D. Madison, and W. Li (LOSS/KAIT; cf. IAUC 8762); and 2006ld by the “Nearby Supernova Factory” collaboration (details on CBET 689).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ke Oct. 19.51 5^{h}52^{m}37.^{s}38 +66^{o}49′00′′.5 18.7 7′′.0 E, 2′′.6 N
2006kf Oct. 21.46 3 41 50.48 + 8 09 25.0 17.4 5′′.5 W, 10′′.6 S
2006ld Oct. 19.3  0 35 27.81 + 2 55 50.7 16.0  —

Additional magnitudes for 2006ke in UGC 3365: Oct. 13.52 UT, [19.0; 20.49, 18.3. SN 2006ke appears to be a subluminous type-Ia supernova around maximum light (details on CBET 684). Additional magnitudes for 2006kf in UGC 2829: Sept. 26.53, [18.8; Oct. 23.45, 17.6. SN 2006kf is a type-Ia supernova before maximum B brightness (details on CBET 691). SN 2006ld in UGC 348 is a type-Ib supernova some 1–2 weeks past maximum (details on CBET 689); nothing was visible at the location of 2006ld on a NEAT image taken on Sept. 20 (limiting mag ≈ 20.0).

COMET C/2006 M4 (SWAN) This comet appears to have undergone an outburst, as indicated by

the following visual total-magnitude and coma-diameter estimates: Oct. 20.75 UT, 5.6, 6′ (A. Diepvens, Balen, Belgium, 20×50 binoculars); 20.79, 5.9, 10′ (J. J. Gonzalez, Asturias, Spain, 7×50 binoculars); 23.75, 5.9, 7′ (R. J. Bouma, Groningen, The Netherlands, 15×80 binoculars); 24.04, 5.4,

– (B. King, Duluth, MN, naked eye); 24.77, 4.0, 7′ (Diepvens); 24.77, 4.4, 8′ (Bouma, 7×50 binoculars); 24.81, 4.5, 8′ (Gonzalez; 1^{o}.8 tail in p.a. 25^{o}; total mag 4.3 via naked eye).

2006 October 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green